# Reassessing Noise Augmentation Methods

### v3

@@ Abstract
In this study, we investigate if noise-augmented training can concurrently improve adversarial robustness in automatic speech recognition (ASR) systems. We conduct a comparative analysis of the adversarial robustness of four different state-of-the-art ASR architectures, where each of the ASR architectures is trained under three different augmentation conditions: one subject to background noise, speed variations, and reverberations, another subject to speed variations only, and a third without any form of data augmentation. The results demonstrate that noise augmentation not only improves model performance on noisy speech but also the model’s robustness to adversarial attacks.
@@
automatic speech recognition, adversarial attacks, adversarial robustness, noise robustness
@@ 1 Introduction
Adversarial attacks, also known as evasion attacks, on machine learning models have garnered significant attention due to their potential to compromise the models’ reliability and security [1]. While early research primarily focused on adversarial attacks on computer vision models, there has been a growing interest in the vulnerability of audio processing models, particularly speech processing systems, to such attacks [2, 3, 4, 5]. Concurrently, efforts to enhance noise robustness have been undertaken to improve the usability of speech processing systems in various environments [6]. However, the relationship between noise robustness and adversarial robustness in the audio domain remains largely unexplored. This paper contributes to filling this gap by investigating whether noise-augmented training can also improve adversarial robustness of speech recognition systems.
@@
Noise robustness, also known as environmental robustness [7] reflects a model’s ability to maintain performance despite exposure to various forms of noise and perturbations. Given that real-world audio data often includes inherent noise and environmental disturbances, ensuring noise robustness is essential for reliable audio processing and analysis [8].
@@
Adversarial robustness, on the other hand, refers to a model’s ability to resist adversarial attacks and maintain accurate predictions when faced with manipulated audio samples. These attacks introduce perturbations that are often imperceptible to humans and can deceive the model into making incorrect predictions [2, 3]. In automatic speech recognition (ASR) systems, adversarial attacks can compromise security and privacy, potentially leading to unauthorized access or incorrect processing of sensitive data. For instance, an attacker could alter an audio signal to mislead an ASR system into executing unintended commands, such as purchasing an unwanted product or navigating to a malicious webpage [4]. Thus, adversarial robustness is critical for model security, enabling the detection and prevention of malicious attacks on audio systems [5].
@@
Understanding the interplay between noise robustness and adversarial robustness in the audio domain is crucial [9]. We hypothesize that ASR models with enhanced resistance to acoustic perturbations may inherently exhibit a degree of robustness to adversarial attacks. To test this hypothesis, we conduct a comprehensive empirical analysis, evaluating the impact of noise-augmented training on the adversarial defense capabilities of various state-of-the-art ASR models. Our study encompasses a range of noise types, in particular background disturbances, and reverberations, and assesses the models’ robustness against attacks in the white-box scenario.
@@ 2 Background and Related Work
This section provides an overview of adversarial attacks and various types of robustness concerning ASR systems. We also discuss the interdependency between adversarial and noise robustness, drawing on literature from the image domain.
@@ 2.1 Adversarial Attacks
Adversarial examples are inputs intentionally modified to induce errors in machine learning models. These perturbations, often imperceptible to humans, can significantly degrade model performance. In speech recognition, adversarial examples exploit ASR system vulnerabilities, resulting in incorrect transcriptions. This study employs a targeted white-box attack to analyze the models’ susceptibility to adversarial attacks.
@@
[2] introduced the first efficient white-box targeted attack against ASR systems, known as the C&W attack. Assuming the attacker has full access to the model, they can create manipulated audio data that transcribes to a specific target phrase. Originally designed for Mozilla’s DeepSpeech model [10], the C&W attack is a gradient-based method that iteratively computes adversarial noise.
The attack minimizes the following loss function with respect to a perturbation $\delta$:
$L(x,\delta,\hat{y})=l_{m}(f(x+\delta),\hat{y})+c\cdot|\delta|_{2}^{2}\enspace,$
where $x$ represents the benign audio signal, $f(\cdot)$ is the ASR model’s function mapping an audio input to the most likely word sequence, $l_{m}$ is the loss function measuring the accuracy of the ASR system’s prediction of the target transcription $\hat{y}$, and $\delta$ is the perturbation constrained to be smaller than a predefined value $\epsilon$. This value is decreased iteratively to minimize distortion. The parameter $c$, which balances the model’s loss against the perturbation, starts small to find an initial adversarial example and then increases to minimize $\delta$.
The resulting adversarial audio files closely resemble the original samples in terms of their dB scale, while leading to a different transcription output by the ASR system. Figure 1 compares the original spectrum of a benign audio clip with its corresponding C&W adversarial counterpart. Note that in line with the definition of adversarial noise, the noise is treated as simply additive.
@@ 2.2 Model Robustness
(a) Original audio clip.
(b) C&W adversarial example.
(c) Adversarial noise signal.
Figure 1: The spectrogram of (a) the original audio signal is compared to (b) the spectrogram of its corresponding C&W adversarial example and (c) the spectrogram of the adversarial noise. Transcription of the original audio signal: “but you know more about that than i do sir”. Transcription of the targeted C&W adversarial example: “yes my dear watson i have solved the mystery”.
@@ 2.2.1 Noise Robustness
Environmental noise is a common challenge in real-world audio processing, requiring ASR systems to handle disturbances like background chatter, traffic noise, and reverberation. General noise, similar to adversarial noise, can cause non-targeted misclassifications, affecting accuracy and user experience. Strategies to enhance noise robustness include noise augmentation, denoising algorithms, and robust feature extraction techniques [11, 12, 13, 14].
Noise-augmented training is widely recognized as a reliable method for improving noise resilience [7]. This technique integrates noise-perturbed data during neural network training by adding specific noise types to clean data. Leading ASR systems, such as SpeechBrain, use noise augmentation in their training protocols, with the selection and weighting of noise types being crucial for effectiveness [15].
While some approaches focus on architectural modifications to inherently enhance noise robustness, such as changes in hidden layers or input data representation [6], we focus on noise-augmented training without altering model architecture or processing pipelines.
(a) Original audio clip.
(b) Noisy original audio clip.
(c) Noise signal.
Figure 2: The spectrogram of the original audio signal (a) is compared to the spectrogram of its corresponding noisy example, which in this case is constructed by adding background noise to the original audio signal (b) and the spectrogram of the noise (c), which is the difference between the original and its noisy counterpart. Transcription of the original audio signal: ”to their sorrow they were soon undeceived”. Predicted text for the noisy benign example: ”to their sorrow they were assumes under seats”.
@@ 2.2.2 Adversarial Robustness
Adversarial robustness concerns a model’s ability to withstand artificially created noise designed to mislead it, as outlined in Section 2.1. Various techniques have been proposed to enhance this robustness, including adversarial training, defensive distillation, gradient masking, regularization techniques, robust optimization, and feature denoising during training [16, 17, 18, 19, 20, 21, 22]. During inference, techniques like input transformations and randomization are employed [23, 24].
Adversarial training involves training the model on both clean and adversarially perturbed samples to enhance its resilience against adversarial manipulations. However, this process is computationally intensive, especially for large datasets. [25] highlighted that while adversarial training is effective for simple datasets, it struggles to scale for high-dimensional data like speech, leaving many blind spots.
This paper evaluates whether noise robustness can enhance adversarial robustness. If successful, this approach could potentially reduce the computational resources needed for generating adversarial examples during training, thereby minimizing blind spots without compromising robustness.
@@ 2.2.3 Correlation between Robustness Notions
Noise robustness enables a model to maintain accuracy amid naturally occurring acoustic disturbances, while adversarial robustness refers to resistance against maliciously engineered inputs. The relationship between these robustness types in the image domain is debated: some studies argue for their independence [26, 27], while others suggest a synergistic relationship [28, 29]. The consensus on their interrelation remains elusive.
Conclusions from image-based research do not directly apply to the audio domain [30]. ASR systems possess unique characteristics, and audio adversarial attacks can exploit phenomena like psychoacoustic hiding, absent in visual contexts [4]. While many studies focus on one type of robustness [3, 4, 7], few have examined both in the audio context [31, 32].
@@ 3.1 ASR Models
We analyze end-to-end ASR systems, which directly transform spoken language into text using a single, unified model, differing from traditional systems with separate stages [33]. We utilize SpeechBrain, an open-source platform for speech technologies [15]. SpeechBrain ASR systems typically consist of three components:
  • Tokenizer: Processes input transcripts to determine corresponding tokens during training and maps output tokens back to text during inference.
  • Acoustic Model: Maps feature vectors derived from the raw input speech into probability distributions over the tokens.
  • Language Model (optional): Enhances prediction accuracy by evaluating the most probable sequences.
@@
We analyze four different architectures from the SpeechBrain toolkit:
@@
For each architecture, we employed three different training regimes:
@@
  • 1 – Baseline (no augmentation): This model serves as a control, trained on a clean dataset without any augmentation to establish standard performance under ideal conditions.
@@
  • 2 – Augmentation with speed variations: Introduces temporal variability in the training data by applying speed perturbations, simulating natural speech tempo variations [41].
@@
  • 3 – Augmentation with speed variations, background noise, and reverberation: Combines speed variations with background noises and reverberations to mimic challenging and realistic acoustic environments [42, 43].
Each model’s architecture and training parameters are detailed in [15].
@@ 3.2 Dataset
We use the LibriSpeech dataset, a widely recognized corpus for ASR research that includes thousands of hours of read English speech from audiobooks [44]. It is often used as a benchmark for evaluating speech recognition models. Our models are trained on the first 100 hours of the clean LibriSpeech dataset. Note that some pre-trained models, like wav2vec, may have been exposed to larger portions of LibriSpeech.

### v4

# Comparative Study on Noise-Augmented Training and its Effect on Adversarial Robustness in ASR Systems

@@ Abstract
In this study, we investigate whether noise-augmented training can concurrently improve adversarial robustness in automatic speech recognition (ASR) systems. We conduct a comparative analysis of the adversarial robustness of four different ASR architectures, each trained under three different augmentation conditions: (1) background noise, speed variations, and reverberations; (2) speed variations only; (3) no data augmentation. We then evaluate the robustness of all resulting models against attacks with white-box or black-box adversarial examples. Our results demonstrate that noise augmentation not only enhances model performance on noisy speech but also improves the model’s robustness to adversarial attacks.
@@
Automatic speech recognition , adversarial examples , adversarial robustness , noise robustness.
journal: Computer Speech and Languages
[…]
@@ 1 Introduction
Adversarial attacks, also known as evasion attacks, on machine learning models have gathered a lot of attention due to their potential to compromise the models’ reliability and security (Szegedy et al., 2014). Initially, research primarily focused on adversarial attacks targeting computer vision models, but there has been a growing interest in the vulnerability of audio processing models, especially speech processing systems, to such attacks. At the same time, efforts to enhance noise robustness have been made to improve the usability of speech processing systems in various environments (Dua et al., 2023). However, the relationship between noise robustness and adversarial robustness in the audio domain remains largely unexplored.
This paper builds upon and extends our previous work (Pizzi et al., 2024) by offering a broader and more rigorous evaluation of the relationship between noise robustness and adversarial robustness in ASR systems. Beyond the earlier study, which focused solely on white-box attacks, the current work incorporates both white-box and black-box adversarial attacks and includes the SI-SDR metric for a perceptual assessment of distortion. These additions enable a more comprehensive and generalizable analysis of how noise-augmented training affects adversarial robustness across diverse threat models and model types.
@@
Noise robustness, also known as environmental robustness Ko et al. (2017), reflects a model’s ability to maintain performance despite exposure to various forms of noise and perturbations. Given that real-world audio data often includes inherent noise and environmental disturbances, ensuring noise robustness is essential for reliable audio processing and analysis (Virtanen et al., 2012).
@@
Adversarial robustness, on the other hand, refers to a model’s ability to resist adversarial attacks and maintain accurate predictions when faced with manipulated audio samples. These attacks introduce perturbations that are often imperceptible to humans but can deceive the model into making incorrect predictions (Carlini and Wagner, 2018; Qin et al., 2019). In automatic speech recognition (ASR) systems, adversarial attacks can compromise security and privacy, potentially leading to unauthorized access or incorrect processing of sensitive data. For instance, an attacker could alter an audio signal to mislead an ASR system into executing commands, such as purchasing an unwanted product or navigating to a malicious webpage (Schönherr et al., 2019). Thus, adversarial robustness is critical for model security, enabling the detection and prevention of malicious attacks on audio systems (Olivier, 2023).
@@
Understanding the interplay between noise robustness and adversarial robustness in the audio domain is crucial (Olivier and Raj, 2023). We hypothesize that ASR models with enhanced resistance to acoustic perturbations may inherently exhibit a degree of robustness to adversarial attacks. To test this hypothesis, we conduct a comprehensive empirical analysis, evaluating the impact of noise-augmented training on the adversarial defense capabilities of various high-performance ASR models. Our study encompasses a range of noise types, such as reverberations and background disturbances, and assesses the models’ robustness against different forms of adversarial evasion attacks, including both white-box and black-box scenarios.
@@ 2 Background and Related Work
This section provides an overview of various types of adversarial attacks against ASR systems as well as of previous results on noise and adversarial robustness of ASR systems. Additionally, we discuss what is known about the interdependence between adversarial and noise robustness, drawing on literature from the image domain.
@@ 2.1 Adversarial Attacks
Adversarial examples are inputs intentionally altered to cause machine learning models to make errors. These perturbations are usually imperceptible to humans and can significantly degrade model performance. In the context of speech recognition, adversarial examples exploit the vulnerabilities of ASR systems, resulting in incorrect transcriptions. Two different classes of adversarial attacks exist, i.e., targeted and untargeted attacks, as described in the following.
2.1.1 Targeted Attacks
@@
Targeted attacks on ARS systems aim at creating manipulated audio data that is transcribed to a specific target phrase chosen by the attacker. These attacks are designed for a white-box setting, which means that the attacker must have full access to the model. Carlini and Wagner (2018) introduced the first efficient white-box targeted attack against ASR systems, known as the C&W attack. Originally designed for Mozilla’s DeepSpeech model (Hannun et al., 2014), the C&W attack computes the adversarial perturbation $\delta$ for a given benign audio sample $x$ by performing gradient descent on the following loss function:
$L(x,\delta,\hat{y})=l_{m}(f(x+\delta),\hat{y})+c\cdot|\delta|_{2}^{2}\kern 5.0pt,$
where $f(\cdot)$ is the function represented by the ASR model mapping the audio input to the most likely transcription, $l_{m}$ is the loss function measuring the accuracy of the ASR system’s prediction w.r.t. the target transcription $\hat{y}$, and $c$ is a hyperparameter regulating the strength of penalization for a larger perturbation. During gradient-based minimization, the perturbation $\delta$ (also referred to as adversarial noise) is constrained to keep smaller than a predefined value $\epsilon$. This $\epsilon$ value is decreased iteratively to minimize the distortion. Moreover, the hyperparameter $c$ is initially set to a small value to simplify finding an adversarial example first. It is then increased to minimize the needed perturbance. The resulting adversarial audio files closely resemble the original samples in terms of their dB scale. However, they lead to a different transcription output by the ASR system. Figure 1 compares the original spectrum of a benign audio signal with its corresponding C&W adversarial counterpart.
(a) Original audio signal. Transcription: “but you know more about that than i do sir”.
(b) C&W adversarial example. Transcription: “yes my dear watson i have solved the mystery”.
(c) Adversarial noise signal $\delta$. Difference between (a) and (b).
Figure 1: The spectrogram of (a) the original audio signal is compared to (b) the spectrogram of its corresponding C&W adversarial example, and (c) the spectrogram of the adversarial noise.
2.1.2 Untargeted Attacks
Untargeted attacks aim to cause a machine learning model to produce an incorrect output without specifying what the incorrect output should be. Unlike targeted attacks, where the adversary has a specific erroneous outcome in mind (e.g., transcribing an audio sample to a specific phrase), untargeted attacks are concerned only with causing the model to fail in its prediction. Untargeted attacks can be particularly useful in scenarios where the goal is to broadly disrupt the functionality of a system or to demonstrate the general vulnerability of a model to adversarial manipulation. These attacks do not require the attacker to have detailed knowledge of the model’s architecture or parameters. In this work, we consider two untargeted black-box attacks for ASR systems (Alzantot et al., 2018; Abdullah et al., 2021). These attacks rely on probing the model with different inputs to generate adversarial examples that lead to incorrect transcriptions without needing access to the internal workings of the model, which makes them particularly interesting under certain attack scenarios. We refer to these attacks as Alzantot and Kenansville attack, respectively.
While Alzantot et al. (2018) use genetic algorithms, an optimization technique that simulates the process of natural selection by repeatedly selecting, combining, and mutating candidate solutions to find the most optimal or fit solution to a problem, Abdullah et al. (2021) present an attack leveraging signal processing methods by systematically removing low-intensity frequency components at each step. They try to remove a minimal number of frequencies that would cause a mismatch between the model’s output and the original label. While the Kenansville attack is actually black-box – they just use the model’s output –, the Alzantot attack can be enhanced with access to the loss function to improve the success rate, which would make it more of a gray-box approach (Olivier and Raj, 2022). Figures 2 and 3 illustrate a comparison between the original spectrum of a benign audio signal and the corresponding adversarial counterparts created with the Alzantot and Kenansville attack, respectively.
(a) Original audio signal. Transcription: “why fade these children of the spring”.
(b) Alzantot adversarial example. Transcription: “why fague these children of the spring”.
(c) Adversarial noise signal $\delta$. Difference between (a) and (b).
Figure 2: The spectrogram of (a) the original audio signal is compared to (b) the spectrogram of its corresponding Alzantot adversarial example, and (c) the spectrogram of the adversarial noise.
@@ 2.2 Model Robustness
(a) Original audio signal. Transcription: “ verse two ”.
(b) Kenansville adversarial example. Transcription: “first too”.
(c) Adversarial noise signal $\delta$. Difference between (a) and (b).
Figure 3: The spectrogram of (a) the original audio signal is compared to (b) the spectrogram of its corresponding Kenansville adversarial example, and (c) the spectrogram of the adversarial noise.
@@ 2.2.1 Noise Robustness
Environmental noise is a common challenge in real-world audio processing, requiring ASR systems to handle disturbances like background chatter, traffic noise, and reverberation. General noise, similar to adversarial noise, can cause non-targeted misclassifications, affecting accuracy and user experience. Strategies to enhance noise robustness include noise augmentation, denoising algorithms, and robust feature extraction techniques (Abe et al., 2015; Stoller et al., 2018; Bagheri and Giacobello, 2019; Alamdari et al., 2021).
Noise-augmented training is widely recognized as a reliable method for improving noise resilience (Ko et al., 2017). This technique integrates noise-perturbed data during neural network training by adding specific noise types to clean data. Modern ASR systems, such as SpeechBrain (Ravanelli et al., 2021), use noise augmentation in their training protocols. The selection and weighting of noise types are crucial for the method’s effectiveness.
While some approaches focus on architectural modifications to inherently enhance noise robustness, such as changes in hidden layers or input data representation (Dua et al., 2023), we focus on noise-augmented training without altering model architecture or processing pipelines.
@@ 2.2.2 Adversarial Robustness
Adversarial robustness concerns a model’s ability to resist adversarial attacks and maintain accurate predictions when faced with manipulated audio samples, as outlined in Section 2.1. Various techniques have been proposed to enhance this robustness, including adversarial training, defensive distillation, gradient masking, regularization techniques, robust optimization, and feature denoising during training (Goodfellow et al., 2015; Madry et al., 2018; Papernot et al., 2016; Tramèr et al., 2018; Jakubovitz and Giryes, 2018; Wong and Kolter, 2018; Xie et al., 2019). During inference, techniques like input transformations and randomization may be employed (Xu et al., 2018; Pizarro et al., 2021).
Adversarial training involves training the model on both clean and adversarially perturbed samples to enhance its robustness against adversarial manipulations. However, this process is considered computationally intensive, especially for large datasets. Zhang et al. (2019) highlight that while adversarial training is effective for simple datasets, it struggles to scale for high-dimensional data like speech, leaving many blind spots.
In this work, we evaluate whether noise robustness can enhance adversarial robustness. If successful, this approach could potentially reduce the computational resources needed for generating adversarial examples during training, thereby minimizing blind spots without compromising robustness.
@@ 2.2.3 Correlation between Robustness Notions
Noise robustness enables a model to maintain accuracy amid naturally occurring acoustic disturbances, whereas adversarial robustness refers to robustness against maliciously engineered inputs. The relationship between these robustness types in the image domain has been debated: some studies argue for their independence (Fawzi et al., 2016; Laugros et al., 2019), while others suggest a synergistic relationship (Gilmer et al., 2019; Li et al., 2019). The consensus on their interrelation remains elusive.
Furthermore, the conclusions from image-based research do not directly apply to the audio domain (Żelasko et al., 2021). ASR systems possess unique characteristics, and audio adversarial attacks can exploit phenomena like psychoacoustic hiding, absent in visual contexts (Schönherr et al., 2019). While many studies focus on one type of robustness (Ko et al., 2017; Qin et al., 2019; Schönherr et al., 2019), few have examined both in the audio context (Olivier and Raj, 2022; Pizarro et al., 2024).
@@ 3.1 ASR Models
We analyze end-to-end ASR systems, which directly transform spoken language into text using a single, unified model, thus differing from traditional systems with separate stages (Prabhavalkar et al., 2024). We utilize SpeechBrain, an open-source platform for speech technologies (Ravanelli et al., 2021). SpeechBrain ASR systems typically consist of three components: (i) during training, a tokenizer breaks the continuous stream of text into discrete normalized units, called tokens (usually characters or subwords) and during inference maps output tokens back to text; (ii) the acoustic model maps feature vectors derived from the raw input speech into probability distributions over the tokens; and (iii) an optional language model that enhances prediction accuracy by evaluating the most likely sequences.
@@
We analyse four different architectures from the SpeechBrain toolkit (using the categories from the SpeechBrain recipes):
  • 1.
    CTC: Uses a unigram tokenizer encoding 31 characters representing the English alphabet and the well-established pre-trained transformer-based model wav2vec 2.0111For the wav2vec 2.0 pre-trained model, see https://huggingface.co/facebook/wav2vec2-large-960h-lv60-self. as the acoustic model (Baevski et al., 2020), a well-established framework for self-supervised learning of speech representations to extract audio features. This model is subsequently fine-tuned by integrating it with a vanilla deep neural network (DNN) that uses a softmax function to learn discrete speech units corresponding to the tokens. The DNN is trained with Connectionist Temporal Classification (CTC) loss (Graves et al., 2006). Since the alignment between the input speech frames with their corresponding textual output is based on CTC, it does not require pre-segmented data, i.e., data divided into smaller chunks. This approach enables the model to learn the alignment explicitly, facilitating the training process for ASR tasks.
  • 2.
    seq2seq 1: Uses a unigram tokenizer, which transforms words into subwords represented in 5000 units. Additionally, it employs an encoder, a decoder, and an attention mechanism between them as the acoustic model. The encoder consists of a combination of convolutional, recurrent, and fully-connected networks, commonly known as CRDNN in the literature (Ravanelli et al., 2021). The decoder employs a long short-term memory network (Hochreiter and Schmidhuber, 1997) coupled with an attention mechanism (Chorowski et al., 2015), initially introduced in the Listen, Attend, and Spell system (Chan et al., 2016). It incorporates a pre-trained language model, trained on a transformer model [^2] For the transformer pre-trained language model, see https://speechbrain.readthedocs.io/en/latest/API/speechbrain.lobes.models.transformer.TransformerLM.html. using the entire LibriSpeech corpus containing approximately 10 million words. This architecture is particularly suited for tasks where the input and output sequences are of variable lengths, making it a popular choice for ASR systems.
  • 3.
    seq2seq 2: Similar to seq2seq 1, but with a tokenizer with only 1000 subword units and a pre-trained recurrent neural network-based language model  [^3] For the RNN pre-trained language model, see https://speechbrain.readthedocs.io/en/latest/API/speechbrain.lobes.models.RNNLM.html. rather than a transformer model.
  • 4.
    transformer: Uses a unigram tokenizer with 5000 sub-word units. It features an acoustic model based on a transformer architecture (Vaswani et al., 2017; Wolf and et al., 2020), which uses attention mechanisms for both encoding and decoding to weight the influence of different parts of the input data. Additionally, it employs the same pre-trained language model as seq2seq 1. This architecture has achieved remarkable success across various domains, including ASR, by effectively capturing the contextual relationships in the data.
@@
The architecture and training parameters of each model are detailed in Ravanelli et al. (2021). For each architecture, we employed three different training regimes:
  • 1.
@@
    No augmentation: This setting serves as a control. The models are trained on a clean dataset without any augmentation to estimate a baseline performance.
  • 2.
@@
    Speed-augmented: This setting introduces temporal variability in the training data by applying speed perturbations that simulate natural speech tempo variations (Ko et al., 2015).
  • 3.
@@
    Fully augmented: For data augmentation in this regime, speed variations are combined with background noises and reverberations to mimic challenging and realistic acoustic environments (Park et al., 2019, 2020).
@@ 3.2 Dataset
We use the LibriSpeech dataset, a widely recognized corpus for ASR research that includes approximately thousand hours of read English speech from audiobooks (Panayotov et al., 2015). It is often used as a benchmark for evaluating speech recognition models. Our models are trained on the first 100 hours of the clean LibriSpeech dataset, which provides a well-balanced platform for conducting controlled, repeatable, and computationally feasible experiments on adversarial robustness across diverse ASR model types. Note that some pre-trained models, like CTC, may have been exposed to larger portions of LibriSpeech.
3.3 Data Augmentation